OpenStudy (anonymous):

I'm having some trouble with limits at infity. The integral is e^(-x)cos(x) dx evaluated from zero to infinity, so the antiderivative is 1/2 e^(-x) ( -cos(x) + sin (x) ). Then since it's from zero to infinity I made it into a limit: lim b--> infinity from 0 to b. Long story short, I don't know how to evaluate it at infinity.

OpenStudy (watchmath):

ok so you have \[ \lim_{b\to \infty} \frac{1}{2}-\frac{1}{2}e^{-b}(-\cos b+\sin b) \] right?

OpenStudy (anonymous):

yeah once you plug in the b and 0, right? \[\frac{ 1 }{ 2 } e^-b (-\cos b + \sin b) - (-1/2)\]